RXTE Helpdesk/FAQ RXTE What's New HEASARC Site Map

RXTE Users' Group Minutes April 2-3, 1996 RXTE

April 2-3, 1996 RXTE Users' Group Meeting

The minutes of XTE Users' Group held on April 2-3 , 1996 at NASA Goddard Space Flight Center are as follows:

XTE User Working Group Meeting Minutes

Tuesday, 2 April, 1996

0830 -  Welcome and Agenda -Schwartz

Dan discussed the charter of this committee and gave his guidelines for
conduct of the meeting.  Hale suggests ascii files instead of
postscript only as a more convenient option for circulation of
viewgraph material.  A mini-poll and discussion concluded that it is
not *.ps  vs ascii per se that is the issue, but the inconvenience in
printing many different *.ps files, or in having very long  *.ps
files.  Remillard said he would set up something easy to use at MIT.
Dan suggests that the e-mail announcement of the availability give us
a listing of file names and lengths, and use some template naming so
that all files can be printed with a single command.

No substantive changes made to the agenda

Dan asks whether HQ wants to consider replacement of inactive current
SWG members.  He suggested that HQ contact them and ask their
intentions or preference for further participation.  Lou replies that we
originally planned to reconsitute the committee once operations began.
There was no specific action assigned on this issue

The date for the next meeting was discussed.  Lou remarked that Guenter
Riegler is planning to have the  Senior Review early this summer.  Jean
thinks this is impossible for XTE, but no sympathy has been offered for
the particular circumstances of XTE.  10-11 Sept, 96 has been chosen as
the next meeting date.  A backup date of 17-18 Oct was identified, to
be used only in case a severe conflict develops.

0835 - HQ Reorganization - Kaluzienski

Lou described the "New improved OSS"  Things are still being defined,
although the new organization was made official 2 to 3 weeks ago. Old
science divisions are gone.  Wes Huntress is  Associate  Administrator
He oversees a board of directors:  Bunner (SEU), Weiler (ASO), Rahe
(SSE), Withbroe(SEC).  Bunner advises on high energy astrophysics and
also gravitational physics.   Weiler has UV sub-UV (search for
origins).  These people make policy (do not administer).  Activities
below the board of directors are divided as follows:  1)Mission and
Payload Development - This will not remain at HQ in the long term.
These activities to be moved to the centers.  The goal is to get down
to 68 people.  2) Adv.Tech & Mission Studies, and 3)  Research Program
Management - Henry Brinton,  Guenter Riegler, David Bohlin, Vernon
Jones, Bredekamp.  SRT grants will be administered from headquarters as
before.  Procurement will be moved to GSFC.  WFF will continue to
administer the rocket SRT program.  SRT programs will look more like
current flight programs from a procurement standpoint.

The impact of the reorganization on XTE should be minimal since most of
the program has been managed at GSFC already.  MO&DA funding is
scheduled to drop after FY97.  As discussed by Riegler, see below,
there is a plan to rephase funding.  They do intend to maintain funds
to support the current AO.

Allen Bunner arrived.  He was asked whether there are plans to
reconstitute the MOWGs.  Allen anticipates that they will be
reconstituted within the SEU sub-disciplines.  Allen posed the question
whether our committee would like to see them at the lower level.  There
was no immediate response.  It was asked where such advice would be
most effective, since the MOWG at Bunner's level would not explicitly
represent X-Ray Astronomy.  Some  sentiment was expressed for a lower
level X-Ray MOWG.

Hale asks whether the current 3 year plan for funding X-Ray MO&DA is
threatened.  The senior review will help determine the level for the
third year.  Note that some of year 2 funding was given up for a "more
rational" third year.  There is some realization that XTE was at a
disadvantage in having to defend itself before it was launched.  Lou
notes a similar problem with the phasing of the next review in that
results are needed in the first 3-4 months of the mission.

HETE is manifested for after November.

Guenter Riegler's presentation, given just before lunch, follows:

Reorganization of the MO&DA program is completed.  No dramatic changes
are expected in the near term.

The date of the senior review has been advanced by a few months to
enable budgetary decisions to be made in a more timely manner.  The
review is now scheduled for summer 96.  The review will address a two
year horizon in detail and the subsequent two years in tentative
fashion.  Guenter listed the nine projects that will be reviewed (ASCA,
GO programs; CGRO; EUVE; HETE; ISO, IPAC support; ISO, US GO program;
ROSAT GO; SAX GO; XTE).  As examples of the overall squeeze in funding
he notes that GRO is planned to receive $5M in its last year compared
to an initial level of $20M.  HETE presently is not scheduled for a
full year of funding.  ISO and XTE are presently scheduled for funding
through FY99.  ROSAT is planned for funding through FY97.  There is no
funding planned for SAX.  So there will obviously be some "competition"
among the missions for resources.  On the other hand, about $20M (out
of $79M) over the next three fiscal years, is not yet earmarked for
specific projects.  According to Guenter, XTE is in the "best shape"
since it will be showing successful operation and actual collection of
data. Funding for XTE was budgeted to start in Sept96, but we didn't
start expending these funds until after the December launch.  There
are plans to rephase part of the unexpended funds (~2.5 to 3.0 $M),
resulting in an MODA profile of 10, 8.3, 7, 6.4 $M from FY96 through
FY99.  The Senior Review will concentrate on Astrophysics MO&DA
programs which have near term problems.  As mentioned earlier, the main
problem is that the overall funding level will decrease sharply in the
next two years.  A possible post 98 funding wedge could develop.  The
pool of funds to be allocated as a result of the review is 50, 36, and
26 $M in FY96, 97, and 98 (excluding HST and AXAF).
Infrastructure/Mission Completion has a relatively constant funding
level.  Individual missions will have the option of continuing
operations vs other ways to spend money (e.g. archiving data).  The
review is now scheduled for the 2nd week of July.  Proposals are due
from the projects in mid-June.  Results will be issued in late August.

Guenter notes that each project will be asked to declare a minimal
funding level, below which it would cease to be viable.  There was some
discussion of how difficult it will be to arrive at such a number for
XTE since we have not het achieved a "normal" equilibrium operating

0922 - Mission Status - Swank

Jean opened with an overview of the PCA status.  Sensitivity is about
.2 mCrab, since fluctuations are seen on that level.  Schwartz asked
about the origin of the fluctuations.  This was the predicted level of
sky noise fluctuations, but they suspect that these are true background
variations since there is some correlation with vetoes.

Jean next gave an overview of the HEXTE status.  She noted that HEXTE
and PCA co-alignments are quite good and represent a real
accomplishment.  The HEXTE Instrument Team has identified a 25-40%
deadtime effect from large energy losses.  This may require longer
observations to reach a specified sensitivity.  HEXTE event processing
does not keep up at 20xCrab rate.

Jean then gave an overview of the ASM status.  Remillard confirmed that
the ASM is operating about 50% of the time, and is turned down to avoid
potential high count rates the remainder of the time.  Calibration on
weaker sources is"coming along"

Overview of the spacecraft:  The telemetry throughput capability is
exceeding specifications (e.g. Cyg X-2 observations used 70Kbps for a
day).  Three problem areas have been identified:  1)  One of two
telemetry transponders loses lock occasionally.  The second unit is
OK.  The fault in the first unit can be bypassed. 2)  The Star trackers
intermittantly lose lock for about 1/3 of the time.  The original
theory of debris in the field of view is not likely to be the real
story.  Spacecraft engineers see no evidence for missing spacecraft
components as culprits.  The spacecraft reacquires stars automatically
after all 5 stars are lost.  3)  Solar arrays show evidence that more
than 59 of the 1737 solar cells are cracked.  They drop out after
warm-up as the spacecraft enters the day portion of orbit.  The
committe asks whether the power system will constrain the mission
life.  The prognosis is that power shouldn't constrain the mission life
if the arrays do not degrade any more.  Operation of the panels at 45
degree solar incidence is not "crucial" according to Kevin Hartnett.

Jean next gave an overview of the science outlook.  Bright sources are
relatively easy to analyze.  AGNs are expected to take longer.  Jean
concerned about how to motivate early observers to get their results
out as soon as possible.  This  generated comments from the audience on
delays in getting data to the observers. This situation would be
discussed more fully later.

Jean jumped ahead to her viewgraph on data analysis.  She notes
problems in implementing FTOOLS.  Fred asks whether the public data is
getting GSFC computer resources for analysis.  There are no plans to
support such analysis using GSFC systems.  It was noted that it is
inefficient to have users download the entirety of the data before
inspecting it.  There was some discussion of how this was planned to
occur.  Nick White recommended extracting one standard EA file, and/or
other standard data products (which are of relatively small size,
examining those, and then going back in a more informed manner to
transfer the data actually needed.

TOOs and RFOs -  There was discussion of the ambiguities of whether a
TOO should trigger an NRA or not.  Frank Marshall notes that we have
not gone to a TOO in less than 7 hours. The problem of keeping PCA warm
complicates rapid response.  Hale wonders whether operations are
settling down and getting more routine.  It was said that there is an
impression that lots of TOO's are being done -- actually, lots of
replanning of EDS modes is taking place, not necessarily just for a
TOO.  The AO-2 makes it clear that one cannot use proprietary HEXTE or
PcA data to trigger a TOO.

BATSE bright source reports have been used, although there is a feeling
that the reports are not being made sufficiently accessible.  The XTE
GOF gets the BATSE bright source list, but is not allowed to distribute
it further.  Dan questions whether the data are proprietary.  Bunner
notes the GRO policy of first claimnant getting rights to new sources.
This ends up being BATSE people in most cases.

It was noted that proposers have a legitimate need for this type of
information.  There appears to be a lack of communication between BATSE
and XTE in this regard.  Dan offers that HQ should pressure BATSE to be
more cooperative to help in management of the XTE observations.   Jean
thinks she can work it out; otherwise, Bunner will help.  Jean notes
that, in general,  XTE observers should alert possible users if they
see appropriate source behavior

The disposition of slew data was discussed.  GOs get the slew data at
the beginning of their observation and part of the slew at the end of
the observation.  Slew data IS public domain, and therefore in
principle is available to anyone to trigger a TOO.  In practice, it is
very difficult to access and use, and the SOC have devoted their own
science research time to making this possible.

PCA problems?  Al Bunner needs a summary before he leaves at noon.  It
was decided to defer discussion of PCA status till after lunch so that
adequate time was available.  We will get him an advance copy of the
meeting minutes on this topic.

1030 -HEXTE Status - Rothschild

HEXTE is generally working well.  There are problems  with one PHA and
with the measurement of dead-time.  Rick notes that the background has
stabilized at about twice the broadband level quoted in AO1.  This is
due to certain background lines (in particular, those associated with
the passive collimators) being more intense that expected.  Rick showed
the particle monitor and anticoincidence shield rates as indicative of
the particle environment.  It was noted that PCA and ASM see softer
particles over a wider portion of the orbit.

The automatic gain control system is working well.  Rick showed plots
of the coincidence alpha rates and the gain state vs time.  He noted
"spikes" in the AGC gain level readout.  Does the gain actually
change?  Check this!

The aperture modulation system has performed as planned.  Rick showed
modulated crab rates.  Lou wonders if any AO1 investigations are
invalidated by the higher background.  The observation time could, in
principle, be extended to compensate for the higher background.  Alan
wonders whether measurement of cylotron features is compromised.  The
appendices for AO2 are being updated to reflect the actual background

Rick described how the counters continue to detect events during
periods for which the high voltage has been reduced for passage through
the SAA.  Telemetry is conserved during these periods by configuring
HEXTE in its idle mode.

Rick showed a plot of the GRO burster light curve indicting the
limiting of event processing due to burden on the processor.  This
occurs only for rates well above the design specifications of the

Rick showed the results of alignment calibrations.  All detectors are
within 5' of the observatory axis.  The energy resolution for all
detectors is degraded by less than 5% of prelaunch value and is well
within the performance specification.

Two major problems have been encountered.  The pulse height analyzer
for one detector has failed, probably due to the failure of an
electrical component.  This detector might still be used as an integral
counter since the phototube and front end electronics still function
correctly.  Al Levine notes that the effective range of the counter
might be modified to perhaps make it more useful.  There is no
explanation for the failure of the PHA component.  Regarding deadtime,
there are two levels of correction needed to the measurement as
generated by the instrument.  A minor correction is needed to account
for small inaccuracies in the gating of the counter for processed
events and discriminator widths.  A more significant problem is the
deadtime due to large energy losses saturating the preamp.  Rick showed
an anti-correlation of good event background rate with XULD rate for
certain extreme conditions of geomagnetic L.  He then showed scope
traces indicating shaping amp and deadtime blanking absent for events
between the 2 MeV and 20 MeV XULD levels.  Finally, he shows lab
measurements of the recovery of the preamp extending beyond the nominal
XULD blanking level.  As a result of these findings, the blanking
window has been extended to 2.5 ms.  This will result in 20-25%
deadtime due to XULD events.  The HEXTE team is investigating some
ideas for mitigating this effect.

Rick listed calibrations to be completed for which data are in hand.
He also showed data from several sources, including a phase shift
measurement of the 1744 pulsar after a burst.

1100 - HEXTE Operations - Gruber

Duane noted that there is a tool available to predict the positions of
the offset background regions for a given source position and time.  He
also noted that considerable effort has gone into assisting user in
selection of HEXTE modes.  He remarked that there has been good
turnaround at the SOC in implementing custom STOL procedures.  He
anticipates some additional engineering observations to investigate
deadtime effects.  HEXTE team is presently  waiting on receipt of
packet data, in particular, some standard products are missing such as
ApID14 aspect data.

The HEXTE response matrix is clean except in the lowest energy channels
and is nearly in form to support general observations.  A background
tool in place and is valid for estimation of background within a16s
instrument data frame.  Extension of the background predicting tool to
shorter time scales is not a high priority.  Background modulation and
subtraction is working well.  As discussed by Rick, the HEXTE team is
still working on the details of the livetime measurement.  The livetime
problem does not affect ones ability to substract background
adequately.  The main effect is on normalization of the subtracted
spectrum.  Jim Lockner claimed that no HEXTE FTOOLS have been delivered
to the GOF.  Dan requested that HEXTE clarify this and report later in
the meeting.

1120 - MODA Reorganization - Riegler

The minutes from this discussion are presented earlier, following Lou
Kaluzienski's presentation

1230 - Lunch

1330 - PCA Performance - Jahoda

Keith discussed various aspects of the PCA performance including dead
time measurement and absolute timing accuracy.  They have verified the
barycenter correction to an absolute accuracy of 0.5 (possibly 1.7) ms
by comparing the x-ray and radio phases of the crab pulsar.  The
boresights have been calibrated and determined to have offsets no
larger than ~5 arcmin, which is within specification.  They have a
model for the collimator response that fits observations assuming a 6
arcmin spreading of the individual tubes.  Keith showed a detailed
readout of the aspect jitter in which the anticipated HEXTE modulation
and ASM stepping perturbations are evident but less than the specified
10 to 30 arcsec level.

Keith showed several examples of rate behavior and spectra indicative
that the PCU's are generally functioning as expected.  There is some
sort of low level ringing effect that occurs in one anode chain (Layer
L2) of one detector (PCU 4) at the lowest of the four possible VLE
settings.  This is not expected to affect the user since the default
window length is longer.  The gain calibrations show good stability
over a 60 day span.  One of the spectra was for a detector that had
exhibited a breakdown anomaly (discussed below) and had recovered.

Keith showed a Crab spectrum of counts residuals that are below the 1%
level below 20 keV.  The deviations appear systematic.

The Collimator Response Tool includes alignment offsets but not
calibrated effective area.  They see a 10-15% discrepancy in absolute
crab rate relative to the "standard" Toor and Seward measurement.  Van
der Klis wonders about the validity of using any particular measurement
of the Crab as a standard.  There was some discussion of the
significance of the 1% discrepancies of the Crab spectrum from a simple
power law.  The PCA response matrix includes quadratic calibration,
escape peak effects and the E**1/2 dependence of the energy

Regarding "the anomaly":  Operation was perfect for 70 days at which
point the 2 coldest detectors (PCU's 3 and 4) exhibited episodic "gentle breakdown".  PCU's 1,2 and 3 have shown no anomalies.

The detectors were then warmed by chosing targets appropriately and the
HV was reduced.  Detectors 4 & 5 seemed to have recovered.  Currently
those two units are turned off and the other units are left on.
Breakdown events are characterized by gain reduction, extra counts in
the affected layers and "no-flag" events.  (No-flag event is an
anomalous event associated with a front end timing glitch and is
generally regarded as understood and not a problem).  The Vx rate
history generally shows a step increase.  The interpretation is that a
gain change shifts events from higher to lower layers.  In every
instance the problem has been corrected by cycling the HV.  Keith
showed a plot of the first moments of R1 vs L1.  For comparison, he
showed a comparable distribution during an event, clearly indicating a
distinctive signature.

They have changed operating procedures.  They now turn off the HV
during SAA (previously cycled only to 1000v).  They have also reduced
the nominal HV by 20v (and by 40v in PCU #5).  This reduced the gain by
~15%.  They are instituting an operational plan to keep the detectors
warmer.  They anticipate a shift in temperature of ~6 C at most.  They
are also programming the EDS to watch for future events and react by
turning the counter off using a process similar to that used for the
ASM.  They are currently studying the expected temperature shifts vs
roll-offset and modified power dissipation distribution in SC.  Jean
notes that such anomalies are more likely at lower operating
temperatures, but she doesn't have an accepted model for the current
behavior.  She speculates that lower temperatures could lead to
build-up of contaminants on the electrodes and eventually lead to
breakdown.  She doesn't think the problem is related to high rates of
normal events.  Will Zhang relates accounts of similar breakdown in
the lab in which the detectors were opened up and buildup was
discovered on the cathode wires in the vicinity of the offending cell.

Implications for users:  They are reluctant to operate the offending
units until they have a reliable method of detecting the breakdown and
immediately shutting them down.  Jean also notes that, in the lab,
similar breakdown episodes appears to be cumulative in the sense that
the time between episodes systematically decreases.  Thus, it is not
prudent to let the units operate in breakdown mode if it can be
avoided.  Rick wonders whether possible side effects of raising PCA
temperatures on HEXTE temperatures have been studied.  Jean suggests
that we might consider using fewer PCUs whenever the science doesn't
require full collecting area.  There was considerable discussion of the
appropriate level of conservatism to use in operating the suspect
dectectors.  Jean doesn't want to take chances.  Hale thinks we might
see a reoccurance soon if it really is a problem.  Will thinks we may
very well lose part of unit 4.  Jean not so sure.  The PCA team is
asking for guidance from this committee as to which operational

A proposal based on the general discussion (phrased by Dan) is to
operate PCA in a contingency mode for a 6 to 8 week period:  1. Turn
off PCUs 4 & 5 (which is the current status).  2. Attempt to warm them
up, via 8 degree roll manuever, spinning up the reaction wheels, and
turn-on the redundant spacecraft computer (not likely to be done since
it is not likely to be effective). 3. Implement an EDS software change
to turn off units upon recognition of breakdown (Ed Morgan thinks that
a SW mod for monitoring the PCA behavior could be ready for test around
12 April). 4. Subsequent to the above, use PCUs 4 & 5 only when above
some minimum temperature, and only when needed to meet the minimum
scientific objectives.  Do the observations that are done well under
these circumstances.  Some objectives would need more time; whereas,
others (e.g. measuring random variability) might not be possible.
Evaluate the operation and replan future use of PCA by ~ June 1.

The committee discussed modifications that are needed in the impending
AO given the PCA situation.  Fred feels that the possible penalty to an
observer's chance of selection should be clear in the AO.  Dan suggests
that the PCA status be clearly documented but that proposers not
advised to address the possibility that they may have only 3
detectors.  The successful proposals be would be reevaluated when the
real situation becomes more clear.  Fred's suggestion:  1) Produce a
bulletin immediately, which advises of the current situation.  Assure
that all who access any proposal information or show interest in
proposing receive this bulletin.  2) Around June 1, we may need to
develop and ask for a suplementary questionaire, which will allow
determination of the feasibility of the proposed observation, and/or
the necessity to request additional time.  It seemed that this should
be developed by F. Marshall in accordance with SOC needs.

We should advise proposers of developments later when they become
known.  We will need to discuss this Weds to clarify the committee's
view on this.

1530 - Spacecraft Status - Hartnett

Anomalies in the power, aspect and telemetry systems were discussed.

1)  Power:  The performance of the solar arrays is below
specification.  The consensus is that cells cracked during launch are
most likely the cause.  Assuming they are cracked randomly, 37 out of
72 strings might be affected.  They do not consider delamination as the
likely cause.  Presently, there is no negative trend in the
performance.  Some of the degradation goes away at lower temperatures.
They are developing procedures to minimize mechanical and thermal
stresses on the arrays.  Al asks about impact on end-of-life
performance.  Currently predicting EOL performance of 800 W relative to
the present load of 650 W assuming no further degradation.

2)  Star Tracker "Loss of Track" (LOT) anomaly:  LOT observed on all 5
guide stars in the day portion of orbit.  Competing theories involve
either debris or earth angle.  NorRad has observed debris in the
vicinity of the spacecraft; however, 16 of 40 LOT events did happen
with no evidence of debris in the position/intensity data.  The
corrective action currently used involves a "TSM".  New ACS code (under
design) will reacquire each guide star as LOT occurs.  They are
considering a proposal to downlink raw pixel data to Ball for
analysis.  The NEAR mission also uses this tracker and is experiencing
similar problems.

3)Transponder:  Unit 2 began to exhibit anomalous behavior after 17
contacts.  The effect is to cause STGT's receiver to lose signal lock.
The anomaly is observed only on unit 2 in "coherent turnaround mode".
TRMM will use the same transponder.  It has flown previously on GRO,
COBE and EUVE.  The anomaly has been localized to the failure of a
component.  A new tracking service allows a work-around that minimizes
the impact of this failure.

Other Ops issues:  There was one case where the ACS Kalman filter
diverged due to lack of timely star updates.  Now are limiting the
slew speed to 6 degrees/minute.  They are recomending a minimum
observation interval of 1000 seconds during next cycle.  There was no
reaction to this recommendation, implying that the committee concurs
with that restriction as announced by Alan Smale in his Feb e-mail.  

Jean asked whether the attitude error is larger than needed due
to modifications to the Kalman filter?  They are evaluating
methods for elevating the PCA temperature.  They might relax the
negative roll bias constraint.  They are also considering increasing
the nominal reaction wheel bias and activation of the redundant OBC as
a heat source.  All anomalies are under control.  They expect answers
on the PCA thermal questions by end of the week.

1600 - SOF Status - Frank Marshall

Staffing:  They plan to hire an additional person to enable Sandy
Antunes to work on scheduling full time.  Frank expects a longer lead
time for schedule changes unless absolutely needed.  There have
typically been  50 targets per week.  The PCA problem has decreased
efficiency.  Instrument teams are augmenting their monitoring
capabilities.  Examples include HEXTE summed background subtracted
spectrum and PCA "layer" light curves.  Mission Monitoring is not
working well.  Frank is not optimistic about improvement soon.  Frank
showed some slew data.  Fred wonders whether slew data will be used to
augment ASM.  This would be an advantage for measuring spectra.  He
suggests that the SOF could do some limited analysis to monitor
selected sources for spectral state, etc.  He is not advocating
setting up specific slew trajectories.  Effects of Instrument
Anomalies:  Have not made any corrections for known anomalies.  Frank
discussed SOF Interfaces with the community.  

1630 - XSDC Status - Bob Patterer

XSDC responsibilities are; 1) Archive raw data from XTE/SOF, 2) Convert
to FITS and Archive, and 3) Distribute to GO's, IT's, MDSS and the
GOF.  The mission Data Staging System provides public data access.
They lost the optical disk archive jukebox. Recovery was delayed by the
government furlough.  SparcStorage Array disk failures have caused
additional problems.  3/4 of AO1 processing is done.  They have
generated 34 days of FITS data.  They have distributed all of Level 0
and most of the FITS data to the Instrument Teams (except HEXTE).
There were some questions about regenerating IOC data and sending it to
the IT's.

1700 -GOF Status - Alan Smale

FITS file and database creation is working well.  FTOOLS version 3.5
has been released.  Public Data are available for 1/19 to 2/1.  He
discussed Data Analysis Software and Calibration Tools (FTOOLS 3.5).
Rick notes that the barycenter correction does not work on HEXTE event
data.  This apparently has a problem on PCA event data also. They
currently are unable to:  1) Subtract background from PCA light curves,
2)  Create a PCA background spectrum, or 3) Apply PCA deadtime
corrections.  Alan Listed some HEXTE issues including:  Analysis of
HEXTE Event Mode data, creation of HEXTE response matrices, subtraction
of background, application of gain and offset corrections and correction
of HEXTE data for collimator response and deadtime.  He discussed Web
Pages and claims these are resulting in good XTE publicity.  Alan showed
but didn't go over changes to the text for AO2.  He asked what changes
would be appropriate for Appendix F?  Rick notes that the HEXTE
components have been updated by Phil Blanco.  Micheil  asks about changes
to the forms that would address mistakes made in the previous AO.  Phil
Blanco has provided some modified forms for the HEXTE modes.

1750 - ASM Status - Al Levine

He is assuming that we have read the relevant ASM status information on
the web pages.  They have encountered four problems:  1) The sun is too
bright, 2) High background regions outside of the SAA, 3) Apparent
episodic diffuse X-ray emission from outside the SSC3 proportional
counter, and 4) Breakdown in SSC's 2 & 3.  Al showed several rate
histories over the orbit demonstrating particle effects in the SSCs.
He discussed analysis of rate vs position on a wire which indicates
that some events are not consistant with x-rays from cosmic sources;
rather they appear like diffuse emission.  Breakdown appears to have
damaged the electrodes by removing some carbon, leaving sharp edges
which are then more prone to further breakdown.  Two anodes in SSC3 and
one anode in SSC2 exhibited these problems and are permanently
inoperative.  All three anodes were cured of breakdown by letting the
breakdown proceed until all carbon was removed.  Thus, 21 of 24 anodes
in the three detectors are now operating normally.  Al listed
corrective measures they are undertaking.  He showed a plot of
excluded regions in Latitude/Longitude space.  These account for ~1/2
of the orbit time.  They have implemented a TSM or "Telemetry and
Statistics Monitor" to control the camera HV.  The data collection duty
cycle is now about 50%.  They were not able to accomplish their planned
objectives in the IOC period.  Real operations started 20 March.  All 3
SSC's are now operating.  They expect to begin releasing derived source
intensities by 1 May.  Dan asked about re-scheduling IOC activities.
There are no plans for IOC observations in the near term.  Van der Klis
asked when ASM will be able to provide triggers to the SOF.  Fred asked
what portion of the sky can be covered.

1815 - ASM calibration- Ron Remillard

Ron noted that, for the past 10 days, source dwells have been obtained
at 36 to 46% efficiency.  He described calibration plans.  They are
analyzing the background of the counters vs position.  They see 3%
variations and are putting these in their model.  They have viewed
ScoX-1 as a strong source to calibrate wire positions.  They see 10%
fewer counts from the Crab than expected.  They also have some
sightings at anomalously low rates.  Ron showed intensities for several
sources.  There are still some systematic problems to solve before the
data can be trusted as a tool for TOO triggering.  He feels they need
one more "iteration".  Ron showed data indicating extensive variability
from GRS 1915+10.  Is this the first ASM TOO?  They estimate their
current detection sensitivity at about 50 mCrab.  Ron described "image
condensations", presumably from Sco X-1 in the edge of the field of
view.  Jean wonders whether their earlier simulation work is still
valid?  That work didn't include some systematic effects that are now

1845 - Break for Dinner - To reconvene Weds AM

0830 - EDS Status - Ed Morgan

He described a fix for the "single-bit bug".  He has since seen
isolated incidents similar to the bug, but they are very rare.  He
notes that a bug in the gain correction has a straightforward
solution.  Ed noted that the spacecraft TSM capability that was used to
deal with ASM breakdown episodes is extremely useful. He is considering
a web page as method for informing users of current available
configurations. They have relaxed the reset frequency since they have
seen Single Event Upsets only as a result of SAA exposure.

0842 - Science Results - Michiel van der Klis

Sco X-1:  He noted minor problems in operations.  Three observations
were obtained over 3 orbits.  They did need to intervene between
observations. They initially specified the Nyquist freq too low and had
problems commanding to higher time resolution.  The command windows
were not as advertised.  They needed to examine power spectra up to
1100 Hz.  The most knowledgable people were not always at SOC to assist
with non-standard computing requirements.  The duty scientists were
apparently not able to support him adequately.  In his third
observation, the source was brighter and they repointed the spacecraft
to moderate the count rates.

Kevin Hartnett advised that it is useful to request additional command
oportunities in advance if you think you might need them.  Should such
capabilites be advertised in the AO?  Keith notes that the vagaries of
the system make this difficult.  Jean/Keith like to keep this informal
Dan requested that we table this till we discuss issues.

Michiel's second Sco observation showed a narrow peak at 1100 Hz.  The
signal is clear, but weak.  In this case, the large collecting area of
the PCA was crucial.  In subsequent analysis, he is able to see a
second peak at 830 Hz.  He notes that "2 LLD " events added to the
sigificance of the power spectrum peak.  He used EDS to recover these
events.  He howed a log/log plot of same data to illustrate the peaks
more clearly.  A shorter VLE window improved observation by extending
frequency range (I DIDN'T GET ALL OF THIS).  The data also show a
normal 6 Hz QPO.  He also sees 45 Hz peak that is probably real.  A
plot of the autocorrelation function shows the effect of the VLE window
which can act as a cutoff filter above a couple of kHz.

1608-52:  This source also shows a 850 Hz peak.  Michiel showed the
temporal behavior of the 1608-52 QPO peak varying from 850 to 900 Hz.

Back to Sco X-1:  Michiel sees both the 830 and 1100 Hz peaks
simultaneously.  He showed a color/color diagram for Sco.  Three
observations clearly show the transition from normal to flaring
branches.  For observations 2 & 3 the flaring branches do not coincide
on the color/color diagram,  probably due to instrumental effects (e.g.
offsetting the collimator may result in reflections that distort the
spectrum).  Todd noted that they may have seen a 1744 burst reflected
into the collimator when offset by more than a degree.  Michiel showed
a Hardness/Intensity diagram indicating subtle systematic effects.

The frequency of the 1100 Hz peak apparently increases with M dot.  The
freqency of the 800 Hz peak marginally shows the same effect.  Could
these peaks be due to Keplerian effects?  Is the 44 Hz peak a beat
frequency effect?

The power spectral plot dramatically shows the dynamic range of PCA .
The 6 Hz and 1100 Hz peaks appear well correlated.  If the 6 Hz peak is
due to radial flow effects, the 1100 Hz peak might indicate a hidden
pulsar signal reflecting and doppler shifting off the radial flow.
(Michiel also has an argument against this hypothesis, but he didn't
give it).  He raises the question as to why the peak is so narrow?

**** - kilo-Hz Osc in Atoll-like bursters - Todd Strohmayer

Todd discussed observations of 4U 1728-34 and 4U1608-52.  He sees a
500-900 Hz QPO peak that is dependent on intensity.  The frequency
variability also correlated with the spectrum in that the RMS variation
increased with energy.  He has not yet seen energy dependent phase
shifts.  He sees multiple components in the power spectra.  He sees a
40 Hz QPO peak in the high intensity state and perhaps also in the low
state.  He sees many weak features in 10-300 Hz range.  The first 4
minute PSD shows a peak at about 800 Hz. Is this real?  He started by
looking at broad-band PSD, then examined the spectral data which appear
to show the same effects.  They see 800 and 40 Hz peaks that probably
vary in frequency.  They also may see a weak feature above 1000 Hz.
There is some kind of complex feature from 100 to 800 Hz.  The QPO
frequency is seen to decrease with intensity as ~2.5 power.  The 800 Hz
peak RMS amplitude increases with energy.  Todd showed high state cross
correlation plots.  It is not clear how the frequency relates to the
intensity in detail.  They also have a low state observation.  The 800
Hz peak looks broader. They see multiple features in the 1728-34 PSD.
Todd showed a color/color diagam for 3 states and EXOSAT PSD's for
comparison.  How are all these effects explained?  He doesn't think it
can be explained by stellar rotation.  If it is Keplarian motion, the
emission would have to originate within 19 km of the neutron star
surface.  The frequecy seems too high for stellar oscillations.  The
amplitudes seem too large.  Could they be due to a beat frequency

**** - Oscillations in 4U160-52 and GROJ1744-28 - Will Zhang

Will shows a srate vs time plot that indicates regular bursting at
intervals of about 120s.  The PSD shows coherent peaks at this
frequency.  A few days later, the burst interval increases to about
150s.  The PSD still shows a coherent peak.  A few hours later the
bursts became less regular correlating with decreasing intensity.  The
bursts are only a 1% effect relative to the overall emission, so the
large PCA area was crucial.  The peaks of the bursts have a BB spectrum
at 0.6 keV.  Will interprets this as Bildsten's rings of fire on a NS
relating to the accretion rate.  The burst interval is much shorter
than the Bildsten model would suggest.

On 1744-28:  Will sees 40 and 60 Hz features.  The intensity changed by
a factor 2, but the frequency hardly changed.  Will also sees a 40 Hz
peak from this source.  Are the bursts related to the 800 Hz QPO?  Is
the 40 Hz peak in 1728-34 of the same mechanism as that in GRO 11744?
Will asks "when will we start seeing usec bursts"(in GRO J1744)?  He
suggests (following Fred Lamb) that we might be looking straight into
the polar cap.  The reduced opacity along this LOS might enable
microsec coherent emission to escape.

**** - Publicity for HEAD mtg - Jean Swank

Jean showed a list of tentative talks for the HEAD session.  Nick White
suggests starting with the exciting scientific topics and not including
the instrumental talks.  Fred emphasizes the importance of "knocking
their socks off".  Lin Cominsky is offering opportunities for press
conferences on Monday and Tuesday.  XTE and GRO would be discussed in
the first one.  Blazar results in second conference

The discussion then turned to where to publish the early XTE results.
It was noted that Nature embargoes results after their submission.
Should there be another press conference at NASA previous to the HEAD
meeting?  Jean feels that the Senior Review should be of highest
priority and that the press conference shouldn't compromise this.  Ann
Cowley suggests a dedicated issue of ApJ letters as a very visable
method of publication.  Can this be done in time?  The hard copy won't
be on the street by the time of the senior review; however, the papers
could have accepted status.  Nick wonders if we can adequately fill a
special issue by then.  It was answered that the entire issue doesn't
have to be XTE articles, only a significant fraction. Nick prefers
emphasizing Nature papers.  Fred thinks that publication in Nature
would take too long. Jean votesfor ApJ.  It was noted that nature
emphasizes cross-disciplinary interest which would be an advantage in
the senior review.  The consensus seems to be that ApJ is the most
practical option.  Jean willtake the lead to get a bunch of XTE papers
into a single issue of ApJ Letters, with all papers due by ~mid-May.
She should attempt to include guest observers in the special issue.

This raised the question of supporting GOs adequately enough to allow
them to have results ready in time.  For example, observers need the
barycentric correction to analyze their data.  Keith thinks its
essentially there, but at less than ultimate accuracy.  Also, that
correction doesn't yet work on event-list data.  Dan suggests that the
SOF should report on the status of this.  Jim Lochner has done this by
e-mail.  Thanks to him for his prompt action.

We returned to the question of the content of the HEAD session.  Dan
feels that Jean should limit the special talks to 5 papers (definitely
not instrument papers).  Fred re-emphasizes the importance of discovery
type topics.  Ann notes that we can defer on defining exactly what all
of the "flashy" results are.  The committee suggests reserving some
time for things that we expect to come out in the next few weeks.  Ron
Remillard feels that we should emphasize the exiciting topics but there
should be some balance with respect to publicizing the functionality of
XTE.  Dan feels that the above scenario, including an expected
contribution on ASM results would be viable.  Still, we must allow Jean
adequate flexibility to reallocate the actual time for various

There is also a GRO session in which XTE results are being presented.
Chuck Dermer is chairing that session and Chryssa Kouveliotou is
organizing it.  The committee discussed where various papers are
currently scheduled.

Discussion turned to the June AAs meeting.  Do we want to showcase the
XTE mission there or at some future AAS meeting?  Ann notes that the
Jan AAS is about the right time frame for having a special XTE
session.  This could be reported at the senior review.  It is too late
right now to do much for the June meeting.  We should try to schedule
an invited talk by a very good speaker at the AAS meeting  Does the
group want to pursue a special meeting?  It is not clear there would be
enough results by that time.  Plan to decide on this issue at the Sept
10-11 meeting.

1115 - Return to Senior Review

I missed some initial discussion of GRO vs XTE competition at the
review.  Lou OFFERED to mediate some agreement as to how to emphasize
the strengths of the two missions vs having US make negative claims
about each other.  Fred notes that XTE received short shrift at the
last review due to poor organization of the review and running out of
time.  How do we minimize the chance that this happens again?  It was
Lou's opinion that this was due to the committee deciding that XTE
didn't really belong in the review, which is now clear that it didn't.

Lou reminds the group that Guenter is asking us to declare the minimum
useful budget for continuing the mission.  There was discussion on the
complications in figuring such a level.  Would code 500 cooperate in
such an exercise?

How do we proceed in generating the proposal for the review?  Fred
suggests that the writing begin in April.  Jean thinks that time is
better spent by XTE scientists doing science during April and deferring
the proposal till May.  Ann suggests writing the proposal from the
beginning to properly convey the changed status of XTE from speculative
to real.  Fred raises the question of how to get the proposal written
in view of Jean's busy schedule?  Dan notes that, as project scientist,
Jean has the official responsibility to submit the proposal, but that
she should tap the relevant people for contributions as soon as
Fred suggests that a detailed outline soon, specifying the text content
and word limits for each subsection, would be an efficient way of
getting things moving.  Fred also notes that word processor standard
would be useful to facilitate managing the inputs from the various
contributors.  Ann suggests using the ApJ standard AASTeX format (used
with LaTeX)  as if the proposal were a scientific paper.  Dan asks Jean
when she will circulate the outline of the proposal.  Jean answers 21

Dan appealed to get copies of viewgraphs, as presented to him or Mike.

1225 - Adjourn

List of Attendees:

Ed Morgan, Ron Remilard, Hale Bradt, Al Levine, Rick Rothschild, Duane
Gruber, Lou Kaluzienski, Michiel van der Klis, Dan Schwartz,Mike
Pelling, Jean Swank, Don Lamb, Keith Jahoda, Paul Hertz, A. Rots, Alan
Smale, GailRohrbach, Frank Marshall, Bob Patterer, Jim Lochner, Tod
Strohmayer, Barry Giles,  Ann Cowley, Jay Sedlar, Kevin Hartnett, Will
Zhang, Nick White, Al Bunner (till noon Tues)

If you have a question about RXTE, please send email to one of our
help desks.

This page is maintained by the RXTE GOF and was last modified on Tuesday, 14-Sep-1999 13:45:33 EDT.